[figure: 9.7 cm]

MULTI-COMPENSATION AND TRI-CRITICAL PHENOMENA

IN PERIODIC AB_{1}AB_{2} SUPERLATTICE WITH S = 1

G. Wiatrowski^{a}, G. Bayreuther^{b}, and F. Bensch^{b} aDepartment of Solid State Physics, University of Łódź, Poland bInstitut für Angewandte Physik, Universität Regensburg, Germany

We have studied an infinite ferrimagnetic AB_{1}AB_{2} superlattice (S = 1) to have direct evidence of the existence of many-compensation and tri-critical effects when a perpendicular single-ion anisotropy is considered. Previous studies have shown the possibility of existence of two (in compositionally bias-modulated TbCo films) or even more compensation points in mixed-spin systems. We present the theory of spin-1 layered structure based on the use of the exact solution for Ising linear chain in external magnetic field (linear cluster approximation-LCA). The free energy calculated from the Bogolyubov inequality allows us to describe: the II-nd as well as the I-st order phase transitions, layered and total magnetization curves. Our results (phase diagrams) indicate the existence of 1 to 4-compensation points for particular values of structure parameters: the magnetic moments, exchange couplings and single-ion anisotropy. Mainly, the role of the latter is verified thoroughly.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : G. Wiatrowski

Address for correspondence : University of Łódź Department of Solid State Physics Pomorska 149/153, PL-90-236 Łódź, Poland

Email address : wiatr@krysia.uni.lodz.pl